# Integrating AI for Human-Centric Breast Cancer Diagnostics: A Multi-Scale and Multi-View Swin Transformer Framework

### v1

@@ III-B Results
The model is trained using a binary cross-entropy with a label smoothing loss function. This loss function helps improve model generalization by penalizing the model less severely on hard examples, effectively smoothing the decision boundary. We used AdamW optimizer, chosen for its ability to combine the benefits of Adam optimization and weight decay regularization, making it particularly suited for this kind of complex model training. Validation is conducted in parallel with training using a separate subset of the data. This approach allows for continuous monitoring of the model’s performance and generalization capabilities, ensuring robustness before deployment. Metrics such as accuracy, sensitivity, specificity, and the area under the ROC curve (AUC) are computed to provide comprehensive insights into the model’s performance across various thresholds and conditions.
Table I illustrates the results obtained from Swin-based back-ends. These results compare two different fusion mechanisms and the performance of the model with missing views. As can be observed, max-pooling fusion seems to be slightly better than the alternative fusion scenarios. This is an interesting observation showing complex fusion is not required, which can contribute to the strength of the feature extraction pipeline. At the same time, it is observed that the performance degradation due to missing views is quite negligible, illustrating the robustness of the proposed pipeline. Finally, Table II illustrates comparison results with recent state-of-the-art [11, 20, 21] further corroborating superiority of the proposed MSMV-Swin framework.
@@
TABLE II: Compression with state-of-the-art.
@@
TMSMV-Swin Framework Convolution Fusion Both-Views $80.32$ $82.27$ $75.07$ $72.32$ $85.88$
@@ IV Conclusion
In this paper, we introduced the MSMV-Swin framework, a hybrid, multi-scale and multi-view Swin Transformer-based architecture aimed at addressing two key challenges in breast cancer diagnosis through mammography, i.e., (i) Reliance on detailed tumor annotations and (ii) Susceptibility to missing views during test time. The MSMV-Swin framework distinguishes itself by utilizing a novel approach that integrates localized and contextual characteristics of tissues surrounding tumors. In other words, by leveraging the Segment Anything Model (SAM), the framework extracts detailed multi-scale information from both CC and MLO views, effectively capturing both localized and contextual features. Additionally, two different feature fusion strategies were employed to ensure robustness by compensating for missing views without relying on synthetic data. Experimental results demonstrate that the MSMV-Swin framework consistently outperforms existing models. Notably, it achieves superior accuracy and robustness even in scenarios where only single-view mammograms are available, with a negligible performance drop compared to dual-view cases. These findings underscore the potential of the MSMV-Swin framework to advance reliable and comprehensive breast cancer diagnostics in real-world clinical settings.

### v2

@@ III-B Results
The model is trained using a binary cross-entropy with a label smoothing loss function. This loss function helps improve model generalization by penalizing the model less severely on complex examples and smoothing the decision boundary. We used AdamW optimizer, chosen for its ability to combine the benefits of Adam optimization and weight decay regularization, making it particularly suited for this kind of complex model training. Validation is conducted in parallel with training using a separate subset of the data. This approach allows continuous monitoring of the model’s performance and generalization capabilities, ensuring robustness before deployment. To evaluate model performance across varying thresholds and conditions comprehensively, we report standard classification metrics, including accuracy, sensitivity, specificity, and the area under the ROC curve (AUC). Table I illustrates the results obtained from Swin-based back-ends. These results compare two fusion mechanisms and the model’s performance with missing views. As can be observed, max-pooling fusion seems to be slightly better than the alternative fusion scenarios. This is an interesting observation, showing that complex fusion is not required, which can contribute to the strength of the feature extraction pipeline. At the same time, it is observed that the performance degradation resulting from missing views is quite negligible, illustrating the robustness of the proposed pipeline. Finally, Table II illustrates the comparison results with recent state-of-the-art [11, 19, 20] further corroborating the superiority of the proposed MSMV-Swin framework.
@@
TABLE II: Comparison with state-of-the-art.
@@
MSMV-Swin Framework Convolution Fusion Both-Views $80.32$ $82.27$ $75.07$ $72.32$ $85.88$
@@ IV Conclusion
In this paper, we introduced the MSMV-Swin framework, a hybrid, multi-scale and multi-view Swin Transformer-based architecture aimed at addressing two key challenges in breast cancer diagnosis through mammography, i.e., (i) Reliance on detailed tumor annotations and (ii) Susceptibility to missing views during test time. The MSMV-Swin framework distinguishes itself by utilizing a novel approach that integrates localized and contextual characteristics of tissues surrounding tumors. In other words, by leveraging SAM, the framework extracts detailed multi-scale information from both CC and MLO views, effectively capturing both localized and contextual features. Additionally, two feature fusion strategies were employed to ensure robustness by compensating for missing views without relying on synthetic data. Experimental results demonstrate that the MSMV-Swin framework consistently outperforms existing models. Notably, it achieves superior accuracy and robustness even in scenarios where only single-view mammograms are available, with a negligible performance drop compared to dual-view cases. These findings underscore the potential of the MSMV-Swin framework to advance reliable and comprehensive breast cancer diagnostics in real-world clinical settings.